Given that the straight line 2x - 3y = 10 passes through the point (t, 4), find the value of t.
Answer by lwsshak3(11628)

Given that the straight line 2x - 3y = 10 passes through the point (t, 4), find the value of t.
..
2x-3y=10
3y=2x-10
y=2x/3-10/3
plug in coordinates (t,4) for (x,y)
4=2t/3-10/3=(2t-10)/3
2t-10=12
2t=22
t=11